Consider the decomposition reaction:           CuCO3.Cu(OH)2   →  2CuO + CO2 + H2O             

       Calculate the theoretical yield of carbon dioxide, CO2, produced by the decomposition reaction if 10.0 g of CuCO3.Cu(OH)2 are consumed. Copper (II) carbonate – copper (II) hydroxide has a molar mass of 221.13 g/mol.

 


1
Expert's answer
2021-02-01T03:47:43-0500

The balanced reaction equation is:

 CuCO3.Cu(OH)2   →  2CuO + CO2 + H2O  .


As one can see from the equation, 1 mole of Copper (II) carbonate – copper (II) hydroxide produces 1 mole of carbon dioxide. The number of the moles of CuCO3.Cu(OH)2 is its mass divided by its molar mass:

"n(CuCO_3\cdot Cu(OH)_2)= \frac{m}{M} = \frac{10}{221.13}=0.04522" mol.

Therefore, the number of the moles of carbon dioxide equals 0.04522 mol also. Its mass will be the number of the moles times its molar mass 44.01 g/mol:

"m(CO_2) = n\cdot M = 0.04522\cdot44.01 = 1.99" g.


Answer: if 10.0 g of CuCO3.Cu(OH)2 are consumed, the theoretical yield of carbon dioxide, CO2, produced by the decomposition reaction is 1.99 g.